# On the Value of Cross-Modal Misalignment in Multimodal Representation Learning

### v6

@@ C.1 Lemmas
$\mathbf{h}_{t}(\mathbf{z}_{t^{\scriptscriptstyle(\theta)}})$ $=\mathbf{h}_{x}(\mathbf{z}_{x})\quad\text{a.s.}\quad\forall\,\mathbf{z}_{x}\sim p_{\mathbf{s}}\,p_{\mathbf{m}_{x}},\;\mathbf{z}_{t^{\scriptscriptstyle(\theta)}}\sim p_{\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}}\mid\mathbf{s},\theta,\rho}\,p_{\mathbf{m}_{t}},$ (12)
which ensures alignment of representations a.s. for any pair $(\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)})$. The substitution of expectations for the image modality in Eq.˜9 is valid because $\mathbf{x}$ follows the pushforward distribution of $\mathbf{z}_{x}$ under the deterministic diffeomorphism $\mathbf{g}_{x}$ (Eq.˜3); a similar argument applies to the text (Eq.˜4).
Equations˜10 and 11 imply that $\mathbf{h}_{x}$ and $\mathbf{h}_{t}$ map the latent variables $\mathbf{z}_{x}$ and $\mathbf{z}_{t^{\scriptscriptstyle(\theta)}}$ onto uniform distributions over $(0,1)^{n}$ (with $n=|\mathbb{I}^{c}_{\rho}|$), since their differential entropy equals to zero.
@@
(Exclusion of Modality-Specific Variables) . We now show that the smooth functions $\mathbf{h}_{x}$ and $\mathbf{h}_{t}$ depend only on latent semantic variables $\mathbf{s}$ (with further exclusion of components in later steps) and not on modality-specific variables $\mathbf{m}_{x}$ or $\mathbf{m}_{t}$.
@@
    $\mathbf{m}_{x}$ is independent of $\mathbf{z}_{t^{\scriptscriptstyle(\theta)}}$: This means changes in $\mathbf{m}_{x}$ do not influence $\mathbf{z}_{t^{\scriptscriptstyle(\theta)}}$. Moreover, since the text generation process $\mathbf{g}_{t^{\scriptscriptstyle(\theta)}}$ is independent of $\mathbf{m}_{x}$, it follows that:
    $\mathbf{h}_{t}(\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}},\mathbf{m}_{t},\mathbf{m}_{x})=\mathbf{h}_{t}(\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}},\mathbf{m}_{t}).$ (13)
@@
    $\mathbf{m}_{t}$ is independent of $\mathbf{z}_{x}$: This means changes in $\mathbf{m}_{t}$ do not influence $\mathbf{z}_{x}$. Similarly, since the image generation process $\mathbf{g}_{x}$ is independent of $\mathbf{m}_{t}$, it follows that:
    $\mathbf{h}_{x}(\mathbf{s},\mathbf{m}_{x},\mathbf{m}_{t})=\mathbf{h}_{x}(\mathbf{s},\mathbf{m}_{x}).$ (14)
@@
$\mathbf{h}_{x}(\mathbf{s},\mathbf{m}_{x})=\mathbf{h}_{t}(\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}},\mathbf{m}_{t},\mathbf{m}_{x}),\quad\text{a.s.}\quad\forall\,\mathbf{z}_{x}\sim p_{\mathbf{s}}\,p_{\mathbf{m}_{x}},\;\mathbf{z}_{t^{\scriptscriptstyle(\theta)}}\sim p_{\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}}\mid\mathbf{s},\theta,\rho}\,p_{\mathbf{m}_{t}}.$ (15)
Consider a small perturbation $\mathbf{m}_{x}\mapsto\mathbf{m}_{x}+\boldsymbol{\varsigma}$, where $\|\boldsymbol{\varsigma}\|$ is arbitrarily small but remains within the open space $\mathcal{M}_{x}$. Since changes in $\mathbf{m}_{x}$ do not influence $\mathbf{h}_{t}$ by statement (c1), we obtain:
$\mathbf{h}_{t}(\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}},\mathbf{m}_{t},\mathbf{m}_{x}+\boldsymbol{\varsigma})=\mathbf{h}_{t}(\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}},\mathbf{m}_{t},\mathbf{m}_{x}).$ (16)
@@
$\mathbf{h}_{x}(\mathbf{s},\mathbf{m}_{x}+\boldsymbol{\varsigma})=\mathbf{h}_{x}(\mathbf{s},\mathbf{m}_{x}).$
By the smoothness of $\mathbf{h}_{x}$ (inherited from the smoothness of $\mathbf{g}_{x}$ and $\mathbf{f}_{x}$), taking $\boldsymbol{\varsigma}\to\mathbf{0}$ gives:
$\frac{\partial\mathbf{h}_{x}}{\partial\mathbf{m}_{x}}=\lim_{\boldsymbol{\varsigma}\to\mathbf{0}}\frac{\mathbf{h}_{x}(\mathbf{s},\mathbf{m}_{x}+\boldsymbol{\varsigma})-\mathbf{h}_{x}(\mathbf{s},\mathbf{m}_{x})}{\boldsymbol{\varsigma}}=0.$
Thus, $\mathbf{h}_{x}$ is independent of $\mathbf{m}_{x}$.
A symmetric argument applies to $\mathbf{m}_{t}$. If $\mathbf{m}_{t}\mapsto\mathbf{m}_{t}+\boldsymbol{\varsigma}$ with $\boldsymbol{\varsigma}$ a small enough perturbation, then by Eq.˜14 in statement (c2), $\mathbf{h}_{x}(\mathbf{s},\mathbf{m}_{x},\mathbf{m}_{t}+\boldsymbol{\varsigma})$ remains unchanged. The invariance condition in Eq.˜12 then forces $\mathbf{h}_{t}(\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}},\mathbf{m}_{t}+\boldsymbol{\varsigma})$ to remain constant w.r.t. $\boldsymbol{\varsigma}$, showing that $\mathbf{h}_{t}$ is independent of $\mathbf{m}_{t}$. Therefore, the learned representations satisfy:
$\mathbf{h}_{x}(\mathbf{z}_{x})$ $=\mathbf{h}_{x}(\mathbf{s}),\quad\text{a.s.}\quad\forall\,\mathbf{z}_{x}\sim p_{\mathbf{s}}\,p_{\mathbf{m}_{x}},$ (17)
$\mathbf{h}_{t}(\mathbf{z}_{t^{\scriptscriptstyle(\theta)}})$ $=\mathbf{h}_{t}(\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}}),\quad\text{a.s.}\quad\forall\,\mathbf{z}_{t^{\scriptscriptstyle(\theta)}}\sim p_{\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}}\mid\mathbf{s},\theta,\rho}\,p_{\mathbf{m}_{t}}.$ (18)
@@
(Exclusion of Omitted Semantic Variables). We now establish that the function $\mathbf{h}_{x}$ is independent of $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}$, where $\mathbb{I}^{c}_{\theta}=\mathbb{I}_{\mathbf{s}}\setminus\mathbb{I}_{\theta}$. In other words, the learned representations do not contain information about the omitted semantic variables that are absent in the corresponding text.
@@
$\mathbf{h}_{x}(\mathbf{s})\;=\;\mathbf{h}_{t}\!\bigl{(}\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}}\bigr{)},\quad\text{a.s.}\quad\forall\,\mathbf{s}\sim p_{\mathbf{s}},\;\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}}\sim p_{\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}}\mid\mathbf{s},\theta,\rho}.$ (19)
Next, we show by contradiction that $\mathbf{h}_{x}$ is independent of $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}$. Suppose, for the sake of contradiction, that there exists a function $\mathbf{h}^{c}_{x}=\mathbf{f}^{c}_{x}\circ\mathbf{g}_{x}$ which depends on at least one component of the omitted semantic variables $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}$. Formally,
$\exists\,l\in\mathbb{I}^{c}_{\theta},\;(\mathbf{s}^{*}_{\mathbb{I}_{\theta}},\mathbf{s}^{*}_{\mathbb{I}^{c}_{\theta}})\in\mathcal{S},\quad\text{such that}\quad\frac{\partial\,\mathbf{h}^{c}_{x}\bigl{(}\mathbf{s}^{*}_{\mathbb{I}_{\theta}},\mathbf{s}^{*}_{\mathbb{I}^{c}_{\theta}}\bigr{)}}{\partial\,s^{*}_{l}}\;\neq\;0.$
By the $C^{1}$ continuity of $\mathbf{h}^{c}_{x}$, guaranteed by the smoothness of both $\mathbf{g}_{x}$ and $\mathbf{f}^{c}_{x}$, and that $\mathcal{S}$ is an open space, it follows that
$\exists\,\eta>0:s_{l}\;\mapsto\;\mathbf{h}^{c}_{x}\!\bigl{(}\mathbf{s}^{*}_{\mathbb{I}_{\theta}},(s_{l},\mathbf{s}^{*}_{\mathbb{I}^{c}_{\theta}\setminus\{l\}})\bigr{)}\quad\text{is strictly monotonic on}\quad(s^{*}_{l}-\eta,\;s^{*}_{l}+\eta),$
@@
$\mathbf{h}^{c}_{x}\!\bigl{(}\mathbf{s}^{*}_{\mathbb{I}_{\theta}},\,(s^{-}_{l},\mathbf{s}^{*}_{\mathbb{I}^{c}_{\theta}\setminus\{l\}})\bigr{)}\;\neq\;\mathbf{h}^{c}_{x}\!\bigl{(}\mathbf{s}^{*}_{\mathbb{I}_{\theta}},\,(s^{+}_{l},\mathbf{s}^{*}_{\mathbb{I}^{c}_{\theta}\setminus\{l\}})\bigr{)}.$ (21)
@@
$\mathbf{h}^{c}_{x}\!\bigl{(}\mathbf{s}^{*}_{\mathbb{I}_{\theta}},\,(s^{-}_{l},\mathbf{s}^{*}_{\mathbb{I}^{c}_{\theta}\setminus\{l\}})\bigr{)}\;=\;\mathbf{h}^{c}_{x}\!\bigl{(}\mathbf{s}^{*}_{\mathbb{I}_{\theta}},\,(s^{+}_{l},\mathbf{s}^{*}_{\mathbb{I}^{c}_{\theta}\setminus\{l\}})\bigr{)}\;=\;\mathbf{h}_{t}\!\bigl{(}\tilde{\mathbf{s}}^{*}_{\mathbb{I}_{\theta}}\bigr{)},$ (22)
@@
Equations˜21 and 22 thus contradict each other. Hence, such a function $\mathbf{h}^{c}_{x}$ cannot exist. Consequently, $\mathbf{h}_{x}$ must be independent of $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}$. Formally,
$\mathbf{h}_{x}(\mathbf{z}_{x})\;=\;\mathbf{h}_{x}(\mathbf{s}_{\mathbb{I}_{\theta}}),\quad\text{a.s.}\quad\forall\,\mathbf{z}_{x}\sim p_{\mathbf{s}}\,p_{\mathbf{m}_{x}}.$ (23)
@@
(i) Justification for the existence of distinct points in Eq.˜20. This follows from the assumption $p_{\mathbf{s}}>0$ a.e. on $\mathcal{S}$ by Asm.˜3.1. From a latent SCM perspective [53, 67], even if a specific semantic component $s_{l}$ in $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}$ is the ancestor node of some other semantic components in $\mathbf{s}_{\mathbb{I}_{\theta}}$, the strict positivity of $p_{\mathbf{s}}$ ensures that the exogenous noise variables are well-defined. Thus, for different values of $s_{l}$, there exist corresponding noise values that keep $\mathbf{s}^{*}_{\mathbb{I}_{\theta}}$ remaining fixed. (ii) What if the unknown causal structure is $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}\to\mathbf{s}_{\mathbb{I}_{\theta}}$? The potential causal influence from $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}$ to $\mathbf{s}_{\mathbb{I}_{\theta}}$ does not resolve the contradiction. Independence here means that, once $\mathbf{s}_{\mathbb{I}_{\theta}}$ is set, there is no direct functional path from $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}$ to the representations $\mathbf{h}_{x}(\mathbf{s}_{\mathbb{I}_{\theta}})$, i.e., the causal influence among them is fully accounted for by the realized value of $\mathbf{s}_{\mathbb{I}_{\theta}}$.
In summary, these arguments show that $\mathbf{h}_{x}$ is genuinely independent of $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}$, even allowing for arbitrary unknown causal interactions among the latent semantic variables.
@@
$\mathbf{h}_{x}(\mathbf{s}_{\mathbb{I}_{\theta}})\;=\;\mathbf{h}_{t}\!\bigl{(}\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}}\bigr{)},\quad\text{a.s.}\quad\forall\,\mathbf{s}\sim p_{\mathbf{s}},\;\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}}\sim p_{\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}}\mid\mathbf{s},\theta,\rho}.$ (24)
Next, we show that $\mathbf{h}_{t}$ must be independent of $\tilde{\mathbf{s}}_{\mathbb{I}_{\rho}}$ by contradiction. Suppose, for a contradiction, that there exist some function $\mathbf{h}^{c}_{t}=\mathbf{f}^{c}_{t}\circ\mathbf{g}_{t^{\scriptscriptstyle(\theta)}}$ which depends on at least on component of the perturbed semantic variables $\tilde{\mathbf{s}}_{\mathbb{I}_{\rho}}$. Formally,
$\exists\,l\in\mathbb{I}_{\rho},\;(\tilde{\mathbf{s}}^{*}_{\mathbb{I}_{\rho}},\tilde{\mathbf{s}}^{*}_{\mathbb{I}^{c}_{\rho}}),\quad\text{such that}\quad\frac{\partial\mathbf{h}^{c}_{t}(\tilde{\mathbf{s}}^{*}_{\mathbb{I}_{\rho}},\tilde{\mathbf{s}}^{*}_{\mathbb{I}^{c}_{\rho}})}{\partial\tilde{s}^{*}_{l}}\neq 0.$
By the $C^{1}$ continuity of $\mathbf{h}^{c}_{t}$ guaranteed by the smoothness of $\mathbf{f}^{c}_{t}$ and $\mathbf{g}_{t^{\scriptscriptstyle(\theta)}}$, for some sufficiently small $\eta>0$, we have the following inequality:
$\mathbf{h}^{c}_{t}\bigl{(}(s^{-}_{l},\tilde{\mathbf{s}}^{*}_{\mathbb{I}_{\rho}\setminus\{l\}}),\tilde{\mathbf{s}}^{*}_{\mathbb{I}^{c}_{\rho}}\bigr{)}\neq\mathbf{h}^{c}_{t}\bigl{(}(s^{+}_{l},\tilde{\mathbf{s}}^{*}_{\mathbb{I}_{\rho}\setminus\{l\}}),\tilde{\mathbf{s}}^{*}_{\mathbb{I}^{c}_{\rho}}\bigr{)},\;\forall\,s^{-}_{l}\in(\tilde{s}^{*}_{l}-\eta,\tilde{s}^{*}_{l}),\,s^{+}_{l}\in(\tilde{s}^{*}_{l},\tilde{s}^{*}_{l}+\eta).$ (25)
@@
$\mathbf{h}^{c}_{t}\bigl{(}(s_{l}^{\prime},\tilde{\mathbf{s}}^{*}_{A\setminus\{l\}}),\mathbf{s}^{*}_{\mathbb{I}_{\theta}\setminus A}\bigr{)}=\mathbf{h}^{c}_{t}\bigl{(}(s_{l}^{\prime\prime},\tilde{\mathbf{s}}^{*}_{A\setminus\{l\}}),\mathbf{s}^{*}_{\mathbb{I}_{\theta}\setminus A}\bigr{)}=\mathbf{h}_{x}(\mathbf{s}^{*}_{A},\mathbf{s}^{*}_{\mathbb{I}_{\theta}\setminus A}).$
This is contradicted by the inequality established in Eq.˜25, which implies that such a $\mathbf{h}^{c}_{t}$ cannot exist. Consequently, any $\mathbf{h}_{t}$ minimizing the loss must be independent of the perturbed semantic variables $\tilde{\mathbf{s}}_{\mathbb{I}_{\rho}}$, i.e.,
$\mathbf{h}_{t}(\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}})=\mathbf{h}_{t}(\mathbf{s}_{\mathbb{I}^{c}_{\rho}}),\quad\text{a.s.}\quad\forall\,\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}}\sim p_{\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}}\mid\mathbf{s},\theta,\rho}.$ (26)

### v7

@@ C.1 Lemmas
$h_{t}(\mathbf{z}_{t^{\scriptscriptstyle(\theta)}})$ $=h_{x}(\mathbf{z}_{x})\quad\text{a.s.}\quad\forall\,\mathbf{z}_{x}\sim p_{\mathbf{s}}\,p_{\mathbf{m}_{x}},\;\mathbf{z}_{t^{\scriptscriptstyle(\theta)}}\sim p_{\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}}\mid\mathbf{s},\theta,\rho}\,p_{\mathbf{m}_{t}},$ (12)
which ensures alignment of representations a.s. for any pair $(\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)})$. The substitution of expectations for the image modality in Eq.˜9 is valid because $\mathbf{x}$ follows the pushforward distribution of $\mathbf{z}_{x}$ under the deterministic diffeomorphism $g_{x}$ (Eq.˜3); a similar argument applies to the text (Eq.˜4).
Equations˜10 and 11 imply that $h_{x}$ and $h_{t}$ map the latent variables $\mathbf{z}_{x}$ and $\mathbf{z}_{t^{\scriptscriptstyle(\theta)}}$ onto uniform distributions over $(0,1)^{n}$ (with $n=|\mathbb{I}^{c}_{\rho}|$), since their differential entropy equals to zero.
@@
(Exclusion of modality-specific variables) . We now show that the smooth functions $h_{x}$ and $h_{t}$ depend only on latent semantic variables $\mathbf{s}$ (with further exclusion of components in later steps) and not on modality-specific variables $\mathbf{m}_{x}$ or $\mathbf{m}_{t}$.
@@
    $\mathbf{m}_{x}$ is independent of $\mathbf{z}_{t^{\scriptscriptstyle(\theta)}}$: This means changes in $\mathbf{m}_{x}$ do not influence $\mathbf{z}_{t^{\scriptscriptstyle(\theta)}}$. Moreover, since the text generation process $g_{t^{\scriptscriptstyle(\theta)}}$ is independent of $\mathbf{m}_{x}$, it follows that:
    $h_{t}(\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}},\mathbf{m}_{t},\mathbf{m}_{x})=h_{t}(\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}},\mathbf{m}_{t}).$ (13)
@@
    $\mathbf{m}_{t}$ is independent of $\mathbf{z}_{x}$: This means changes in $\mathbf{m}_{t}$ do not influence $\mathbf{z}_{x}$. Similarly, since the image generation process $g_{x}$ is independent of $\mathbf{m}_{t}$, it follows that:
    $h_{x}(\mathbf{s},\mathbf{m}_{x},\mathbf{m}_{t})=h_{x}(\mathbf{s},\mathbf{m}_{x}).$ (14)
@@
$h_{x}(\mathbf{s},\mathbf{m}_{x})=h_{t}(\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}},\mathbf{m}_{t},\mathbf{m}_{x}),\quad\text{a.s.}\quad\forall\,\mathbf{z}_{x}\sim p_{\mathbf{s}}\,p_{\mathbf{m}_{x}},\;\mathbf{z}_{t^{\scriptscriptstyle(\theta)}}\sim p_{\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}}\mid\mathbf{s},\theta,\rho}\,p_{\mathbf{m}_{t}}.$ (15)
Consider a small perturbation $\mathbf{m}_{x}\mapsto\mathbf{m}_{x}+\boldsymbol{\varsigma}$, where $\|\boldsymbol{\varsigma}\|$ is arbitrarily small but remains within the open space $\mathcal{M}_{x}$. Since changes in $\mathbf{m}_{x}$ do not influence $h_{t}$ by statement (c1), we obtain:
$h_{t}(\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}},\mathbf{m}_{t},\mathbf{m}_{x}+\boldsymbol{\varsigma})=h_{t}(\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}},\mathbf{m}_{t},\mathbf{m}_{x}).$ (16)
@@
$h_{x}(\mathbf{s},\mathbf{m}_{x}+\boldsymbol{\varsigma})=h_{x}(\mathbf{s},\mathbf{m}_{x}).$
By the smoothness of $h_{x}$ (inherited from the smoothness of $g_{x}$ and $f_{x}$), taking $\boldsymbol{\varsigma}\to\mathbf{0}$ gives:
$\frac{\partial h_{x}}{\partial\mathbf{m}_{x}}=\lim_{\boldsymbol{\varsigma}\to\mathbf{0}}\frac{h_{x}(\mathbf{s},\mathbf{m}_{x}+\boldsymbol{\varsigma})-h_{x}(\mathbf{s},\mathbf{m}_{x})}{\boldsymbol{\varsigma}}=0.$
Thus, $h_{x}$ is independent of $\mathbf{m}_{x}$.
A symmetric argument applies to $\mathbf{m}_{t}$. If $\mathbf{m}_{t}\mapsto\mathbf{m}_{t}+\boldsymbol{\varsigma}$ with $\boldsymbol{\varsigma}$ a small enough perturbation, then by Eq.˜14 in statement (c2), $h_{x}(\mathbf{s},\mathbf{m}_{x},\mathbf{m}_{t}+\boldsymbol{\varsigma})$ remains unchanged. The invariance condition in Eq.˜12 then forces $h_{t}(\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}},\mathbf{m}_{t}+\boldsymbol{\varsigma})$ to remain constant w.r.t. $\boldsymbol{\varsigma}$, showing that $h_{t}$ is independent of $\mathbf{m}_{t}$. Therefore, the learned representations satisfy:
$h_{x}(\mathbf{z}_{x})$ $=h_{x}(\mathbf{s}),\quad\text{a.s.}\quad\forall\,\mathbf{z}_{x}\sim p_{\mathbf{s}}\,p_{\mathbf{m}_{x}},$ (17)
$h_{t}(\mathbf{z}_{t^{\scriptscriptstyle(\theta)}})$ $=h_{t}(\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}}),\quad\text{a.s.}\quad\forall\,\mathbf{z}_{t^{\scriptscriptstyle(\theta)}}\sim p_{\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}}\mid\mathbf{s},\theta,\rho}\,p_{\mathbf{m}_{t}}.$ (18)
@@
(Exclusion of omitted semantic variables). We now establish that the function $h_{x}$ is independent of $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}$, where $\mathbb{I}^{c}_{\theta}=\mathbb{I}_{\mathbf{s}}\setminus\mathbb{I}_{\theta}$. In other words, the learned representations do not contain information about the omitted semantic variables that are absent in the corresponding text.
@@
$h_{x}(\mathbf{s})\;=\;h_{t}\!\bigl(\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}}\bigr),\quad\text{a.s.}\quad\forall\,\mathbf{s}\sim p_{\mathbf{s}},\;\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}}\sim p_{\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}}\mid\mathbf{s},\theta,\rho}.$ (19)
Next, we show by contradiction that $h_{x}$ is independent of $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}$. Suppose, for the sake of contradiction, that there exists a function $h^{c}_{x}=f^{c}_{x}\circ g_{x}$ which depends on at least one component of the omitted semantic variables $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}$. Formally,
$\exists\,l\in\mathbb{I}^{c}_{\theta},\;(\mathbf{s}^{*}_{\mathbb{I}_{\theta}},\mathbf{s}^{*}_{\mathbb{I}^{c}_{\theta}})\in\mathcal{S},\quad\text{such that}\quad\frac{\partial\,h^{c}_{x}\bigl(\mathbf{s}^{*}_{\mathbb{I}_{\theta}},\mathbf{s}^{*}_{\mathbb{I}^{c}_{\theta}}\bigr)}{\partial\,s^{*}_{l}}\;\neq\;0.$
By the $C^{1}$ continuity of $h^{c}_{x}$, guaranteed by the smoothness of both $g_{x}$ and $f^{c}_{x}$, and that $\mathcal{S}$ is an open space, it follows that
$\exists\,\eta>0:s_{l}\;\mapsto\;h^{c}_{x}\!\bigl(\mathbf{s}^{*}_{\mathbb{I}_{\theta}},(s_{l},\mathbf{s}^{*}_{\mathbb{I}^{c}_{\theta}\setminus\{l\}})\bigr)\quad\text{is strictly monotonic on}\quad(s^{*}_{l}-\eta,\;s^{*}_{l}+\eta),$
@@
$h^{c}_{x}\!\bigl(\mathbf{s}^{*}_{\mathbb{I}_{\theta}},\,(s^{-}_{l},\mathbf{s}^{*}_{\mathbb{I}^{c}_{\theta}\setminus\{l\}})\bigr)\;\neq\;h^{c}_{x}\!\bigl(\mathbf{s}^{*}_{\mathbb{I}_{\theta}},\,(s^{+}_{l},\mathbf{s}^{*}_{\mathbb{I}^{c}_{\theta}\setminus\{l\}})\bigr).$ (21)
@@
$h^{c}_{x}\!\bigl(\mathbf{s}^{*}_{\mathbb{I}_{\theta}},\,(s^{-}_{l},\mathbf{s}^{*}_{\mathbb{I}^{c}_{\theta}\setminus\{l\}})\bigr)\;=\;h^{c}_{x}\!\bigl(\mathbf{s}^{*}_{\mathbb{I}_{\theta}},\,(s^{+}_{l},\mathbf{s}^{*}_{\mathbb{I}^{c}_{\theta}\setminus\{l\}})\bigr)\;=\;h_{t}\!\bigl(\tilde{\mathbf{s}}^{*}_{\mathbb{I}_{\theta}}\bigr),$ (22)
@@
Equations˜21 and 22 thus contradict each other. Hence, such a function $h^{c}_{x}$ cannot exist. Consequently, $h_{x}$ must be independent of $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}$. Formally,
$h_{x}(\mathbf{z}_{x})\;=\;h_{x}(\mathbf{s}_{\mathbb{I}_{\theta}}),\quad\text{a.s.}\quad\forall\,\mathbf{z}_{x}\sim p_{\mathbf{s}}\,p_{\mathbf{m}_{x}}.$ (23)
@@
(i) Justification for the existence of distinct points in Eq.˜20. This follows from the assumption $p_{\mathbf{s}}>0$ a.e. on $\mathcal{S}$ by Asm.˜3.1. From a latent SCM perspective [53, 67], even if a specific semantic component $s_{l}$ in $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}$ is the ancestor node of some other semantic components in $\mathbf{s}_{\mathbb{I}_{\theta}}$, the strict positivity of $p_{\mathbf{s}}$ ensures that the exogenous noise variables are well-defined. Thus, for different values of $s_{l}$, there exist corresponding noise values that keep $\mathbf{s}^{*}_{\mathbb{I}_{\theta}}$ remaining fixed. (ii) What if the unknown causal structure is $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}\to\mathbf{s}_{\mathbb{I}_{\theta}}$? The potential causal influence from $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}$ to $\mathbf{s}_{\mathbb{I}_{\theta}}$ does not resolve the contradiction. Independence here means that, once $\mathbf{s}_{\mathbb{I}_{\theta}}$ is set, there is no direct functional path from $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}$ to the representations $h_{x}(\mathbf{s}_{\mathbb{I}_{\theta}})$, i.e., the causal influence among them is fully accounted for by the realized value of $\mathbf{s}_{\mathbb{I}_{\theta}}$.
In summary, these arguments show that $h_{x}$ is genuinely independent of $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}$, even allowing for arbitrary unknown causal interactions among the latent semantic variables.
@@
$h_{x}(\mathbf{s}_{\mathbb{I}_{\theta}})\;=\;h_{t}\!\bigl(\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}}\bigr),\quad\text{a.s.}\quad\forall\,\mathbf{s}\sim p_{\mathbf{s}},\;\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}}\sim p_{\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}}\mid\mathbf{s},\theta,\rho}.$ (24)
Next, we show that $h_{t}$ must be independent of $\tilde{\mathbf{s}}_{\mathbb{I}_{\rho}}$ by contradiction. Suppose, for a contradiction, that there exist some function $h^{c}_{t}=f^{c}_{t}\circ g_{t^{\scriptscriptstyle(\theta)}}$ which depends on at least on component of the perturbed semantic variables $\tilde{\mathbf{s}}_{\mathbb{I}_{\rho}}$. Formally,
$\exists\,l\in\mathbb{I}_{\rho},\;(\tilde{\mathbf{s}}^{*}_{\mathbb{I}_{\rho}},\tilde{\mathbf{s}}^{*}_{\mathbb{I}^{c}_{\rho}}),\quad\text{such that}\quad\frac{\partial h^{c}_{t}(\tilde{\mathbf{s}}^{*}_{\mathbb{I}_{\rho}},\tilde{\mathbf{s}}^{*}_{\mathbb{I}^{c}_{\rho}})}{\partial\tilde{s}^{*}_{l}}\neq 0.$
By the $C^{1}$ continuity of $h^{c}_{t}$ guaranteed by the smoothness of $f^{c}_{t}$ and $g_{t^{\scriptscriptstyle(\theta)}}$, for some sufficiently small $\eta>0$, we have the following inequality:
$h^{c}_{t}\bigl((s^{-}_{l},\tilde{\mathbf{s}}^{*}_{\mathbb{I}_{\rho}\setminus\{l\}}),\tilde{\mathbf{s}}^{*}_{\mathbb{I}^{c}_{\rho}}\bigr)\neq h^{c}_{t}\bigl((s^{+}_{l},\tilde{\mathbf{s}}^{*}_{\mathbb{I}_{\rho}\setminus\{l\}}),\tilde{\mathbf{s}}^{*}_{\mathbb{I}^{c}_{\rho}}\bigr),\;\forall\,s^{-}_{l}\in(\tilde{s}^{*}_{l}-\eta,\tilde{s}^{*}_{l}),\,s^{+}_{l}\in(\tilde{s}^{*}_{l},\tilde{s}^{*}_{l}+\eta).$ (25)
@@
$h^{c}_{t}\bigl((s_{l}^{\prime},\tilde{\mathbf{s}}^{*}_{A\setminus\{l\}}),\mathbf{s}^{*}_{\mathbb{I}_{\theta}\setminus A}\bigr)=h^{c}_{t}\bigl((s_{l}^{\prime\prime},\tilde{\mathbf{s}}^{*}_{A\setminus\{l\}}),\mathbf{s}^{*}_{\mathbb{I}_{\theta}\setminus A}\bigr)=h_{x}(\mathbf{s}^{*}_{A},\mathbf{s}^{*}_{\mathbb{I}_{\theta}\setminus A}).$
This is contradicted by the inequality established in Eq.˜25, which implies that such a $h^{c}_{t}$ cannot exist. Consequently, any $h_{t}$ minimizing the loss must be independent of the perturbed semantic variables $\tilde{\mathbf{s}}_{\mathbb{I}_{\rho}}$, i.e.,
$h_{t}(\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}})=h_{t}(\mathbf{s}_{\mathbb{I}^{c}_{\rho}}),\quad\text{a.s.}\quad\forall\,\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}}\sim p_{\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}}\mid\mathbf{s},\theta,\rho}.$ (26)